OpenStudy (anonymous):

7(sqrt)13 + (sqrt)6 - 9(sqrt)13

OpenStudy (anonymous):

like terms are \[7\sqrt{13}\] and \[-9\sqrt{13}\] comine to get \[-2\sqrt{13}\]

OpenStudy (anonymous):

-2sqrt13+sqrt6

OpenStudy (nenadmatematika):

correct :D